# FLASH: Latent-Aware Semi-Autoregressive Speculative Decoding for Multimodal Tasks

### v2

@@ Abstract
Large language and multimodal models (LLMs and LMMs) exhibit strong inference capabilities but are often limited by slow decoding speeds. This challenge is especially acute in LMMs, where visual inputs typically comprise more tokens with lower information density than text — an issue exacerbated by recent trends toward finer-grained visual tokenizations to boost performance. Speculative decoding has been effective in accelerating LLM inference by using a smaller draft model to generate candidate tokens, which are then selectively verified by the target model, improving speed without sacrificing output quality. While this strategy has been extended to LMMs, existing methods largely overlook the unique properties of visual inputs and depend solely on text-based draft models. In this work, we propose FLASH (Fast Latent-Aware Semi-Autoregressive Heuristics), a speculative decoding framework designed specifically for LMMs, which leverages two key properties of multimodal data to design the draft model. First, to address redundancy in visual tokens, we propose a lightweight latent-aware token compression mechanism. Second, recognizing that visual objects often co-occur within a scene, we employ a semi-autoregressive decoding strategy to generate multiple tokens per forward pass. These innovations accelerate draft decoding while maintaining high acceptance rates, resulting in faster overall inference. Experiments show that FLASH significantly outperforms prior speculative decoding approaches in both unimodal and multimodal settings, achieving up to 2.68$\times$ speed-up on video captioning and 2.55$\times$ on visual instruction tuning tasks compared to the original LMM. Our code is available [here].
@@ 1 Introduction
Large multimodal models (LMMs)  Achiam et al. (2023); Team et al. (2024); Yang et al. (2024); Liu et al. (2023) have made significant progress on tasks like video captioning and visual question answering by leveraging synergistic relationships between data types. Recent studies show that processing more input tokens during inference improves contextual understanding Muennighoff et al. (2025); Liu et al. (2025). Consequently, LMMs increasingly adopt finer-grained patch division strategies, which significantly expand the token count. This issue is particularly severe for video inputs, as the large number of the video frames leads to a significant expansion of the visual input scale. While larger models and longer contexts improve accuracy and flexibility, they also raise deployment challenges due to hardware constraints and increased computational costs.
@@
To address the efficiency issue, recent works propose token compression to remove some visual tokens Zhang et al. (2025); Wen et al. (2025). Since the visual tokens contain superfluous or non-essential information, pruning these tokens makes shorter input contexts and thus accelerating the generation. However, this simplification comes with potential drawbacks. For example, it can be difficult to accurately identify which tokens to prune, as seemingly non-critical tokens may actually encode latent task-specific cues Wen et al. (2025). Therefore, while token compression offers efficiency gains, it may damage the performance by oversimplifying the visual input, especially in fine-grained multimodal tasks.
@@
Speculative decoding offers a promising solution for accelerating inference without sacrificing performance. As an effective strategy for speeding up decoding in Large Language Models (LLMs) Bachmann et al. (2025); Li et al. (2024b); Cai et al. (2024); Fu et al. (2024), it employs a lightweight draft model to quickly generate candidate token sequences, which are then verified in parallel by the target model through a single forward pass. Crucially, by applying specific acceptance-rejection criteria, the accepted tokens can be regarded as samples from the target model’s distribution, thereby preserving output quality. This fail-safe property ensures that speculative decoding maintains the accuracy of the target model while significantly reducing inference latency. However, the overall speed-up is closely linked to the acceptance rate of the candidates generated by the draft model. This relationship establishes a critical trade-off in the design of draft models: overly simplistic architectures, while computationally efficient, risk producing low-quality candidates that are frequently rejected during verification, diminishing overall latency gains.
@@
Given that LMMs frequently leverage LLMs as their decoders, prior work Gagrani et al. (2024) has attempted to directly transplant speculative decoding techniques from LLMs to the multimodal settings. However, this method trains a separate draft model using only textual inputs, which introduces additional computational overhead and ignores the visual modality. To overcome this drawback, we propose to incorporate information from both visual and textual modalities into the draft model while minimizing the computational costs, thereby enabling more efficient and effective multimodal inference. Recently, Eagle Li et al. (2024b) proposed a simple speculative decoding framework for LLMs to enable efficient inference, in which a lightweight autoregressive head serves as the draft model and takes the second-to-top layer features from the target model as input. However, in contrast to LLMs, LMMs contain numerous redundant visual tokens, making a naive extension of this framework to multimodal input computationally expensive. To address this limitation, we propose FLASH (Fast Latent-Aware Semi-Autoregressive Heuristics), a novel method for efficient multimodal speculative decoding that achieves a favorable trade-off between inference speed and draft quality.
@@
In this work, we leverage two distinctive properties of multimodal data to enhance the efficiency of FLASH: visual token redundancy and vision object co-occurrence. Based on these properties, we design two novel components in the draft model: visual token compression and semi-autoregressive head. Unlike LLMs, LMMs often process a large number of redundant visual tokens, which significantly slow down inference during speculative decoding. To mitigate this, we compress the visual tokens based on the hidden state features, which accelerates draft generation while minimizing the loss of semantic information. Speculative decoding speeds up autoregressive generation using a lightweight draft model to predict the output of a heavier target model. However, since the draft model remains autoregressive, the overall speed-up is limited. In contrast to textual input, visual input inherently exhibits spatial co-occurrence rather than left-to-right causal relationships. Image patches are arranged to reflect their spatial positions, not sequential dependencies. As a result, when describing multiple visual regions simultaneously, models often rely on fixed collocation patterns such as “in front of” or “on the table”, which do not require strict autoregressive ordering. This observation motivates our adoption of a semi-autoregressive decoding strategy, which better preserves spatial relationships while maintaining generation efficiency.
@@
By combining latent-aware visual token compression with semi-autoregressive decoding, FLASH achieves faster inference and maintains high draft quality. We evaluate FLASH on video captioning and visual instruction tuning tasks, using LLaVA Shang et al. (2024) and QwenVL Bai et al. (2025) as target models. Experimental results demonstrate that the two components of FLASH, visual token compression and semi-autoregressive decoding, provide distinct yet synergistic advantages on different tasks. In video captioning, where a large number of visual tokens are involved, visual token compression is particularly effective, while in instruction tuning, which typically involve fewer visual tokens but longer textual inputs, the semi-autoregressive decoding contributes more significantly to efficiency gains. Overall, FLASH achieves average speed-up gains of $24.4\%$ (0.41×0.41$\times$0.41 ×) and $41.5\%$ (0.62×0.62$\times$0.62 ×) on video captioning and visual instruction tuning tasks, compared to the previous methods that rely solely on text tokens in multimodal speculative decoding.
The main contributions are concluded as follows:
@@
(1) To the best of our knowledge, we propose FLASH, the first speculative decoding framework designed for LMMs, which effectively exploits the characteristics of multimodal inputs.
@@
(2) Observing that visual information is often redundant and descriptions of visual content typically appear as phrases, we introduce visual token compression and semi-autoregressive generation to accelerate the draft inference.
(3) Experimental results demonstrate that FLASH significantly speeds up the drafting process in LMMs without causing substantial degradation in draft quality.
@@ 2 Related Works
Speculative decoding. Since the introduction of speculative sampling Leviathan et al. (2023); Chen et al. (2023), the strategy of using light-weight models to generate drafts, with large models performing parallel verification, has been widely adopted to accelerate inference across various Large Language Models (LLMs) Xia et al. (2024b); Gao et al. (2025). However, selecting an appropriate draft model is challenging, as it is difficult to make the predicted distribution consistent with that of the target model due to differences in model size or architecture Bachmann et al. (2025). It has been suggested that knowledge distillation applied to the target model can produce a compact model with a higher reception rate Zhou et al. (2023). Unlike training a draft model independently, self-speculative decoding introduces a way to reuse components of the target model Liu et al. (2024a); Elhoushi et al. (2024); Xia et al. (2024a). Following this thought, Eagle Li et al. (2024b, c) introduces an autoregressive head on the second-to-top feature extracted by the target model to predict candidate tokens. Beyond increasing the acceptance rate of draft tokens, enhancing the drafting efficiency further contributes to achieving a higher speed-up ratio in speculative decoding. Medusa applies n-head architecture, where each head predicts one corresponding token Cai et al. (2024). Lookahead Decoding leverages an n-gram pool generated through Jacobi iterations, enabling the model to accept multi-token prefixes Fu et al. (2024). The aforementioned methods accelerate inference through speculative decoding in LLMs. However, when applied to large multimodal models (LMMs), a key challenge arises: effectively integrating both visual and textual modalities while maintaining efficiency and accuracy during inference.
@@
Token Compression. Prior studies have shown that incorporating additional inputs can help correct erroneous outputs and enhance model performance Liu et al. (2025). However, as Large Multimodal Models (LMMs) continue to scale, improving inference efficiency has become increasingly important Liu et al. (2023, 2024b); Team et al. (2024). A key challenge lies in the high computational cost caused by the large number of visual tokens. To address this, reducing the number of input tokens without sacrificing essential information is considered a crucial strategy. To reduce redundancy, similar vision tokens can be merged based on their similarity Shang et al. (2024); Li et al. (2024a). Image token reduction can also be achieved through a Q-Former Li et al. (2023) to extract of visual concepts Yang et al. (2024); Chen et al. (2024a). However, evidence from recent research indicates that Q-Former leads to some degree of visual information loss Yao et al. (2024); Fan et al. (2024). Based on the observation that early-layer visual tokens contain more critical information, LLaVA-mini integrates visual features into text tokens through pre-fusion before these layers Zhang et al. (2025). DART prioritizes the removal of duplicate tokens over the selection of important tokens Wen et al. (2025).
@@
Figure 1: Illustration of FLASH. The target model is depicted on the left, while the draft model is shown on the right. In this example, the number of the total input tokens $M$ is equal to $N+2$, as the input includes $N$ visual tokens and $2$ textual tokens. The draft model takes the second-to-top layer features $F_{V}$ as the visual input. These features are first compressed by a visual compressor, producing the compressed visual features $\hat{F}_{V}$. Along with textual feature and embeddings $\hat{F}_{T}$, they are fed into a semi-autoregressive head to generate the next $K$ tokens in parallel. These candidate drafts are highlighted with red-bordered boxes.
3 FLASH
@@ 3.1 Preliminaries
In this paper, our Large Multimodal Models (LMMs) process two data modalities, images and texts. As shown in Figure 1, the image is encoded by a vision encoder and projected into visual embeddings $V=\{V_{1},V_{2}...,V_{N}\}$, while the text is encoded into textual embeddings $E=\{E_{N+1},E_{N+2},...,E_{M}\}$, where $N$ denotes the number of visual tokens, and $M$ represents the total number of input tokens, encompassing both visual and textual components. Subsequently, these embeddings are concatenated and processed through Transformer layers to produce the second-to-top-layer feature $F=\{F_{1},F_{2},...,F_{M}\}$. Finally, an LM head maps the last feature $F_{M}$ to a probability distribution over the output vocabulary space, from which the next token $T_{0}$ is sampled.
In speculative decoding, the LMM acts as the target model for output validation, while a lightweight draft model proposes $K$ candidate tokens ${T}^{\prime}_{1},...,{T}^{\prime}_{K}$. During validation, the target model computes the probabilities of these $K$ candidate tokens in parallel within a single forward pass. The acceptance probability of each candidate token is defined as the ratio of its probability under the target model to its probability under the draft model. Subsequently, the tokens are evaluated after obtaining the acceptance probability of each candidate token. For the $i$-th token $T^{\prime}_{i}$, if accepted, it is retained as part of the output. If rejected, the token and all subsequent candidate tokens, $T^{\prime}_{i}$ to $T^{\prime}_{K}$ are discarded, and the draft model is triggered to generate the next $K$ tokens starting from the position $i$.
Given the single-step inference time $\mathcal{M}_{T}$ and $\mathcal{M}_{D}$ for the target and draft model respectively, draft model’s inference time $\mathcal{M}_{D}$ is much less than that of the target model $\mathcal{M}_{T}$, due to the lightweight architecture of the draft model. Besides, the verification of $K$ candidate tokens by the target model is performed in parallel, so the time required to verify all $K$ tokens can be approximated by the single-step time $\mathcal{M}_{T}$. For speculative decoding, the total time includes the autoregressive draft model generating $K$ tokens and the subsequent verification of theses tokens by the target model, which can be expressed as:
@@
When the target model is selected, $\mathcal{M}_{T}$ remains constant. Therefore, to improve the speed-up ratio, it is crucial to reduce the draft generation time $K\cdot\mathcal{M}_{D}$, while maintaining a high draft quality to ensure a large value of $i$. To achieve this, we propose a novel method named FLASH, which effectively balances inference speed and draft quality by leveraging the characteristics of multimodal inputs. Following the previous LLM speculative decoding method Eagle Li et al. (2024b), we construct the textual input $\hat{F}_{T}$ by concatenating the second-to-top layer textual feature $F_{T}$ and the token embeddings $E$. For the visual modality, FLASH introduces a visual token compression module that reduces computational overhead while preserving key visual semantics. To further accelerate decoding, FLASH incorporates a semi-autoregressive head that predicts the next $K$ tokens in parallel, significantly reducing the draft generation time from $K\cdot\mathcal{M}_{D}$ to approximately $\mathcal{M}_{D}$.

### v3

# SpecFLASH: A Latent-Guided Semi-autoregressive Speculative Decoding Framework for Efficient Multimodal Generation

@@ Abstract
Large language models and large multimodal models (LLMs and LMMs) deliver strong generative performance but suffer from slow decoding, a problem that becomes more severe when handling visual inputs, whose sequences typically contain many more tokens with lower information density than text. Speculative decoding accelerates LLM inference by letting a compact draft model propose candidate tokens that are selectively accepted by a larger target model, achieving speed-up without degrading quality. However, existing multimodal speculative decoding approaches largely ignore the structural characteristics of visual representations and usually rely on text-only draft models. In this paper, we introduce SpecFLASH, a speculative decoding framework tailored to LMMs that explicitly exploits multimodal structure when designing the draft model. We first mitigate redundancy in visual token sequences with a lightweight, latent-guided token compression module that compacts visual features while preserving semantics, and then leverage the co-occurrence and local correlations of visual entities via a semi-autoregressive decoding scheme that predicts multiple tokens in a single forward pass. Extensive experiments demonstrate that SpecFLASH consistently surpasses prior speculative decoding baselines, achieving up to $2.68\times$ speed-up on video captioning and $2.55\times$ on visual instruction tuning, relative to the original LMM. Our code is available [here].
Machine Learning, ICML
@@ 1 Introduction
Large multimodal models (LMMs)  (Achiam et al., 2023; Team et al., 2024; Yang et al., 2024; Liu et al., 2023) have made significant progress on tasks like video captioning and visual question answering by leveraging synergistic relationships between data types. Recent studies show that processing more input tokens during inference improves contextual understanding (Muennighoff et al., 2025; Liu et al., 2025). As a result, LMMs increasingly adopt finer-grained patch division strategies, which significantly expand the token count. This issue is particularly severe for video inputs, as the large number of the video frames leads to a significant expansion of the visual input scale. While larger models and longer contexts improve accuracy and flexibility, they also raise deployment challenges due to hardware constraints and increased computational costs.
@@
To address the efficiency issue, recent works propose token compression to remove some visual tokens (Zhang et al., 2025; Wen et al., 2025). Since the visual tokens contain superfluous or non-essential information, pruning these tokens makes shorter input contexts and thus accelerating the generation. However, this simplification comes with potential drawbacks. For example, it can be difficult to accurately identify which tokens to prune, as seemingly non-critical tokens may actually encode latent task-specific cues (Wen et al., 2025). Therefore, while token compression offers efficiency gains, it may damage the performance by oversimplifying the visual input, especially in fine-grained multimodal tasks.
@@
Speculative decoding offers a promising solution for accelerating inference without sacrificing performance. As an effective strategy for speeding up decoding in Large Language Models (LLMs) (Bachmann et al., 2025; Li et al., 2024c; Cai et al., 2024; Fu et al., 2024), it employs a lightweight draft model to quickly generate candidate token sequences, which are then verified in parallel by the target model through a single forward pass. Crucially, by applying specific acceptance-rejection criteria, the accepted tokens can be regarded as samples from the target model’s distribution, thereby preserving output quality. This fail-safe property ensures that speculative decoding maintains the accuracy of the target model while significantly reducing inference latency. However, the overall speed-up is closely linked to the acceptance rate of the candidates generated by the draft model. This relationship establishes a critical trade-off in the design of draft models: overly simplistic architectures, while computationally efficient, risk producing low-quality candidates that are frequently rejected during verification, diminishing overall latency gains.
@@
Given that LMMs frequently leverage LLMs as their decoders, prior work (Gagrani et al., 2024) has attempted to directly transplant speculative decoding techniques from LLMs to the multimodal settings. However, this method trains a separate draft model using only textual inputs, which introduces additional computational overhead and ignores the visual modality. To overcome this drawback, early efforts (Ganesan et al., 2025; Hu et al., 2025) have explored incorporating both visual and textual information into the draft model. However, in contrast to LLMs, LMMs contain numerous redundant visual tokens, making a naive extension of this framework to multimodal input computationally expensive. To address this limitation, we propose SpecFLASH (Speculative Decoding with Fast Latent-Aware Semi-autoregressive Heuristics), a novel method for efficient multimodal speculative decoding that achieves a favorable trade-off between inference speed and draft quality.
@@
In this work, we leverage two distinctive properties of multimodal data to enhance the efficiency of SpecFLASH: visual token redundancy and vision object co-occurrence. Based on these properties, we design two novel components in the draft model: visual token compression and semi-autoregressive head. Unlike LLMs, LMMs often process a large number of redundant visual tokens, which significantly slow down inference during speculative decoding. To mitigate this, we compress the visual tokens based on the hidden state features, which accelerates draft generation while minimizing the loss of semantic information. Speculative decoding speeds up autoregressive generation using a lightweight draft model to predict the output of a heavier target model. However, since the draft model remains autoregressive, the overall speed-up is limited. We observe that although relational phrases (e.g., “on the table” and “in front of”) are not unique to LMMs, they become substantially more constrained and deterministic when conditioned on an image. In text-only generation, such relations are determined purely by discourse intent and can vary widely given the same prefix. In contrast, in vision–language generation, once the model identifies salient objects and their spatial configuration (e.g., a cup on a table), the corresponding relational description strongly co-occurs and is far more predictable. This observation motivates our adoption of a semi-autoregressive decoding strategy, which better preserves spatial relationships while maintaining generation efficiency.
@@
By combining latent-aware visual token compression with semi-autoregressive decoding, SpecFLASH achieves faster inference and maintains high draft quality. We evaluate SpecFLASH on video captioning and visual instruction tuning tasks, using LLaVA (Shang et al., 2024) and QwenVL (Bai et al., 2025) as target models. Experimental results demonstrate that the two components of SpecFLASH, visual token compression and semi-autoregressive decoding, provide distinct yet synergistic advantages on different tasks. In video captioning, where a large number of visual tokens are involved, visual token compression is particularly effective, while in instruction tuning, which typically involves fewer visual tokens but longer textual inputs, the semi-autoregressive decoding contributes more significantly to efficiency gains. Overall, SpecFLASH achieves average speed-up gains of $24.4\%$ and $41.5\%$ on video captioning and visual instruction tuning tasks, compared to the previous methods that rely solely on text tokens.
We make the following contributions:
@@
(1) We propose SpecFLASH, a novel speculative decoding framework for LMMs that effectively exploits the characteristics of multimodal inputs.
@@
(2) By observing that visual information is often redundant and descriptions of visual content typically appear as short phrases, we introduce visual token compression and semi-autoregressive generation to accelerate draft inference.
(3) Experiments show that SpecFLASH achieves substantial acceleration without degradation in output quality.
@@ 2 Related Works
Speculative decoding. Since the introduction of speculative sampling (Leviathan et al., 2023; Chen et al., 2023), the strategy of using light-weight models to generate drafts, with large models performing parallel verification, has been widely adopted to accelerate inference across various LLMs (Xia et al., 2024b; Gao et al., 2025). However, selecting an appropriate draft model is challenging, as it is difficult to make the predicted distribution consistent with that of the target model due to differences in model size or architecture (Bachmann et al., 2025). It has been suggested that knowledge distillation applied to the target model can produce a compact model with a higher reception rate (Zhou et al., 2023). Unlike training a draft model independently, self-speculative decoding introduces a way to reuse components of the target model (Liu et al., 2024a; Elhoushi et al., 2024; Xia et al., 2024a). Following this idea, Eagle (Li et al., 2024c, b) introduces an autoregressive head on the second-to-top feature extracted by the target model to predict candidate tokens. Beyond increasing the acceptance rate of draft tokens, enhancing the drafting efficiency further contributes to achieving a higher speed-up ratio in speculative decoding. Medusa applies n-head architecture, where each head predicts one corresponding token (Cai et al., 2024). Lookahead Decoding leverages an n-gram pool generated through Jacobi iterations, enabling the model to accept multi-token prefixes (Fu et al., 2024). Speculative decoding has also been extended beyond LLMs to LMMs (Ganesan et al., 2025; Hu et al., 2025; Wang et al., 2025). MASSV (Ganesan et al., 2025) adapts speculative decoding to multimodal settings via self-distillation, and Dream (Hu et al., 2025) enhances draft quality through a specialized cross-attention mechanism on visual and textual features. Despite these advances, applying speculative decoding to LMMs introduces a fundamental challenge: how to effectively integrate both visual and textual modalities while preserving efficiency and accuracy during inference.
@@
Token Compression. Prior studies have shown that incorporating additional inputs can help correct erroneous outputs and enhance model performance (Liu et al., 2025). However, as Large Multimodal Models (LMMs) continue to scale, improving inference efficiency has become increasingly important (Liu et al., 2023, 2024b; Team et al., 2024). A key challenge lies in the high computational cost caused by the large number of visual tokens. To address this, reducing the number of input tokens without sacrificing essential information is considered a crucial strategy. To reduce redundancy, visual tokens can be merged based on their similarity (Shang et al., 2024; Li et al., 2024a). Image token reduction can also be achieved through a Q-Former (Li et al., 2023) to extract visual concepts (Yang et al., 2024; Chen et al., 2024b). However, evidence from recent research indicates that Q-Former leads to some degree of visual information loss (Yao et al., 2024; Fan et al., 2024). Based on the observation that early-layer visual tokens contain more critical information, LLaVA-mini integrates visual features into text tokens through pre-fusion before these layers (Zhang et al., 2025). DART prioritizes the removal of duplicate tokens over the selection of important tokens (Wen et al., 2025).
@@
Figure 1: Illustration of SpecFLASH. The target model is depicted on the left, while the draft model is shown on the right. In this example, the number of the total input tokens $M$ is equal to $N+2$, as the input includes NN visual tokens and $2$ textual tokens. The draft model takes the second-to-top layer features $F_{V}$ as the visual input. These features are first compressed by a visual compressor, producing the compressed visual features $\hat{F}_{V}$. Along with textual feature and embeddings $\hat{F}_{T}$, they are fed into a semi-autoregressive head to generate the next $K$ tokens in parallel. These candidate drafts are highlighted with red-bordered boxes.
3 SpecFLASH
@@ 3.1 Preliminaries
In this paper, Large Multimodal Model (LMM) processes two data modalities, images and texts. As shown in Figure 1, the image is encoded by a vision encoder and projected into visual embeddings $V=\{V_{1},V_{2}...,V_{N}\}$, while the text is encoded into textual embeddings $E=\{E_{N+1},E_{N+2},...,E_{M}\}$, where $N$ denotes the number of visual tokens, and $M$ represents the total number of input tokens, encompassing both visual and textual components. Subsequently, these embeddings are concatenated and processed through Transformer layers to produce the second-to-top-layer feature $F=\{F_{1},F_{2},...,F_{M}\}$. Finally, an LM head maps the last feature $F_{M}$ to a probability distribution over the output vocabulary space, from which the next token $T_{0}$ is sampled.
In speculative decoding, the LMM acts as the target model for output validation, while a lightweight draft model proposes $K$ candidate tokens ${T}^{\prime}_{1},...,{T}^{\prime}_{K}$. During validation, the target model computes the probabilities of these $K$ candidate tokens in parallel within a single forward pass. The acceptance probability of each candidate token is defined as the ratio of its probability under the target model to its probability under the draft model. Subsequently, the tokens are evaluated after obtaining the acceptance probability of each candidate token. For the $i$-th token $T^{\prime}_{i}$, if accepted, it is retained as part of the output. If rejected, the token and all subsequent candidate tokens, $T^{\prime}_{i}$ to $T^{\prime}_{K}$ are discarded, the output token at position $i$ is sampled from the residual distribution constructed from the target and draft probabilities (Leviathan et al., 2023).
Given the single-step inference time $\mathcal{M}_{T}$ and $\mathcal{M}_{D}$ for the target and draft model respectively, draft model’s inference time $\mathcal{M}_{D}$ is much less than that of the target model $\mathcal{M}_{T}$, due to the lightweight architecture of the draft model. Additionally, the verification of $K$ candidate tokens by the target model is performed in parallel, so the time required to verify all $K$ tokens can be approximated by the single-step time $\mathcal{M}_{T}$. For speculative decoding, the total time includes the autoregressive draft model generating $K$ tokens and the subsequent verification by the target model, which can be expressed as:
@@
When the target model is selected, $\mathcal{M}_{T}$ remains constant. Therefore, to improve the speed-up ratio, it is crucial to reduce the draft generation time $K\cdot\mathcal{M}_{D}$, while maintaining a high draft quality to ensure a large value of $i$. To achieve this, we propose a novel method named SpecFLASH, which effectively balances inference speed and draft quality by leveraging the characteristics of multimodal inputs. Following the previous LLM speculative decoding method Eagle (Li et al., 2024c), we construct the textual input $\hat{F}_{T}$ by concatenating the second-to-top layer textual feature $F_{T}$ and the token embeddings $E$. For the visual modality, SpecFLASH introduces a visual token compression module that reduces computational overhead while preserving key visual semantics. To further accelerate decoding, SpecFLASH incorporates a semi-autoregressive head that predicts the next $K$ tokens in parallel, significantly reducing the draft generation time from $K\cdot\mathcal{M}_{D}$ to approximately $\mathcal{M}_{D}$.